OpenStudy (anonymous):

Evaluate Sin^-1(aquare roote of 3/2)

OpenStudy (anonymous):

let \[\sin^{-1} \sqrt{\frac{ 3 }{ 2 }}=\] theta

OpenStudy (anonymous):

@nahal_amani are u there?

OpenStudy (anonymous):

yes im here

OpenStudy (anonymous):

Could you use the link @esamalaa ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sin \theta=\sqrt{\frac{ 3 }{ 2 }}\]

OpenStudy (anonymous):

what is the angle that its sin is squaroot 3/2

OpenStudy (anonymous):

theta=60 degrees

OpenStudy (anonymous):

or 240 degrees

OpenStudy (anonymous):

wait let me check

OpenStudy (anonymous):

its 120 degrees

OpenStudy (anonymous):

yes this also can be right

OpenStudy (anonymous):

and sorry 240 degrees is false

OpenStudy (anonymous):

it's 60 and 120 degrees

OpenStudy (anonymous):

Thank you